OpenStudy (anonymous):

in an antiderivative, does the power rule applies to all functions with an exponent on the variable?

OpenStudy (lgbasallote):

hmm do you mean \[\huge \int u^n du = \frac{u^{n+1}}{n+1}?\]

OpenStudy (lgbasallote):

and yeah it applies :)

OpenStudy (zarkon):

\[n\neq -1\]

OpenStudy (lgbasallote):

^when n = -1 you use \[\int u^{-1} = \ln u\]

OpenStudy (lgbasallote):

\[\int u^{-1} du = \ln u\]

OpenStudy (anonymous):

thank you!!!!

OpenStudy (zarkon):

to be really anal we should probably write it as \(\ln|u|\)